OpenStudy (anonymous):

Brian has 3 more books than Erika. Jasmine has triple the number of books that Brian has. Altogether Brian, Erika and Jasmine have 22 books. How many books does Jasmine have?

OpenStudy (campbell_st):

let erika have x books brian has x + 3 jasmine has 3(x +3) so together then have x + x + 3 + 3(x + 3) = 22 solve for x

OpenStudy (anonymous):

let no of book with Erica be x then Brian will have x +3 book and Jasmine will have 3 (3+x) { as it given "Jasmine has triple the number of books that Brian has "} so now total no of book = \[x+ (x+3) +{3 \times (x+3)}=22\]

OpenStudy (anonymous):

5x+12=22 5x=10 x=2 no of book Erica has is 2